OpenStudy (bbb911):

___(y - 5)^2___ y^2 - 25

OpenStudy (anonymous):

Do you want to simplify it ?

OpenStudy (bbb911):

multiply and simplify

OpenStudy (anonymous):

First factor denominator using this formula If we have a^2 - b^2 = (a - b)*(a + b) Here at denominator we have y^2 - 25 = y^2 - 5^2 = (y-5)*(y+5)

OpenStudy (anonymous):

SO at denominator we have got (y -5) * (y + 5) So we get (y - 5)^2 / (y -5) * (y + 5) Then just try to cancel common factor or same terms from numerator & denominator So what do you get ?

OpenStudy (anonymous):

Are u there ?

OpenStudy (bbb911):

this might take a while

OpenStudy (anonymous):



OpenStudy (anonymous):

Have u understood ?

OpenStudy (bbb911):

a little bit im taking online classes and im trying to understand the problems, i am not that good in math

OpenStudy (anonymous):

Can you tell me which step u r not understanding ?

OpenStudy (bbb911):

the elimination of the number like what you did on the top

OpenStudy (bbb911):

oh is it because they are the same so you eliminate them both right?

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

Is it clear now ?

OpenStudy (bbb911):

yes now i get it thank you

OpenStudy (anonymous):

u r welcome :)

OpenStudy (bbb911):

so the answer is ( y - 5) ( y+ 5)

OpenStudy (anonymous):

(y -5) / (y+5) ------------- --------------

OpenStudy (bbb911):

yay! :)

OpenStudy (anonymous):

:)